# Coherence-Driven Multimodal Safety Dialogue with Active Learning for Embodied Agents

### v1

@@ Abstract.
When assisting people in daily tasks, robots need to accurately interpret visual cues and respond effectively in diverse safety-critical situations, such as sharp objects on the floor. In this context, we present M-CoDAL, a multimodal-dialogue system specifically designed for embodied agents to better understand and communicate in safety-critical situations. The system leverages discourse coherence relations to enhance its contextual understanding and communication abilities. To train this system, we introduce a novel clustering-based active learning mechanism that utilizes an external Large Language Model (LLM) to identify informative instances. Our approach is evaluated using a newly created multimodal dataset comprising 1K safety violations extracted from 2K Reddit images. These violations are annotated using a Large Multimodal Model (LMM) and verified by human annotators. Results with this dataset demonstrate that our approach improves resolution of safety situations, user sentiment, as well as safety of the conversation. Next, we deploy our dialogue system on a Hello Robot Stretch robot and conduct a within-subject user study with real-world participants. In the study, participants role-play two safety scenarios with different levels of severity with the robot and receive interventions from our model and a baseline system powered by OpenAI’s ChatGPT. The study results corroborate and extend the findings from automated evaluation, showing that our proposed system is more persuasive and competent in a real-world embodied agent setting.
@@
Embodied Agents, Multimodal Safety, Active Learning, Coherence Theory
@@ 1. Introduction
Embodied agents, such as assistive robots designed for household tasks and daily assistance, should perceive and interpret their surroundings and engage in conversations with human users when noticing an unsafe scenario. The advent of powerful multimodal models (202, 2023; Liu et al., 2023) can allow robots to interpret and respond to cues from both visual and conversational inputs in real-world scenarios. However, ensuring user safety in these interactions requires systems that understand and adapt to the conversational context and remain robust across diverse situations. To address this dual challenge, we integrate theories of coherence relations to enhance contextual understanding and employ clustering-based active learning to train a multimodal dialogue system, M-CoDAL, specifically designed for embodied agents to provide safety advice to users.
@@
M-CoDAL uses discourse coherence relations to enrich its contextual understanding. Theories of coherence relations originate in understanding and analyzing inferential links to support text interpretation and has subsequently been extended to cross-modal settings (Alikhani et al., 2023). Coherence relations can situate an ongoing scene in the arc of a narrative (Cohn, 2013) or enrich interpretation of communicative actions across modalities (Lascarides and Stone, 2009). For instance, the coherence relation Cause can help interpret why paying attention to a safety violation is important in an image captured by a robot given a certain Condition relation. These relations can then facilitate dialogue between the human user and the robot. By default, a robot equipped with pretrained models would not explicitly model such inferential links (Alikhani et al., 2023). We hypothesize that enabling robots to parse coherence relations of safety violation in captured images, and to use these relations to guide their responses, will result in embodied agents that better understand contexts of safety and engage in safer conversations with humans.
To train M-CoDAL for embodied agents with integration of coherence relations, we adopt clustering-based active learning (Settles, 2009). Active learning offers advantage over standard fine-tuning methods by focusing resources toward informative instances. When combined with clustering, active learning has been shown to yield more representative models (Hassan and Alikhani, 2023a). Thus, we hypothesize that applying clustering-based active learning would lead to better coverage of safety scenarios and a safer multimodal dialogue system. Active learning, however, has mostly been addressed in the context of classification tasks and has remained challenging for generative tasks (Perlitz et al., 2023) such as conversational response generation due to difficulty in estimating model uncertainty in a large output space. We make the first inroads in active learning for autonomous agents with conversational capacity by integrating an external LLM that quantifies informativeness of an instance based on a composite score of safety of conversation, resolution, and user sentiment. Further, we distill knowledge from another external LLM to reduce necessary human efforts in active learning. Our work is also one of the first to deploy and evaluate active learning in practice for human-robot interactions.
@@
Finally, we deploy our multimodal dialogue system on a Hello Robot Stretch robot and conduct a within-subject user study where real-world participants interact with our robot in household scenarios regarding potential safety violations. 8 participants interact with the robot operating either with GPT-4o or our proposed system, M-CoDAL in a low and high severity fake unsafe scenario. The study results corroborate findings of automated evaluation by showing that participants found our proposed system M-CoDAL to be more persuasive in both conditions. Participants also find M-CoDAL to be more competent in low-serverity situations. Our qualitative analysis of interviews with the participants also reveals that the participants find M-CoDAL to be more attentive to safety situation and can help the user to be more aware of safety situations. Thus, the key contributions of this paper are:
@@ 2. Related Work
There has been a growing interest in development of multimodal dialogue systems in recent years (Sun et al., 2022a; Lee et al., 2023; Sicilia et al., 2023), especially with the advent of large multimodal models such as GPT-4V (202, 2023) and LLaVA (Liu et al., 2023). Discussions of safety however, has primarily focused on textual domain (Sun et al., 2022b; Weidinger et al., 2022; Dinan et al., 2022; Atwell et al., 2022) and safety of embodied multimodal agents remain relatively unexplored despite its importance. To our knowledge, our work is the first public work to process visual cues of safety in multimodal dialogues with an embodied agent.
@@
Coherence relations have been proposed as a possible method for controlling generative models and have been shown to aid tasks such as extractive and abstractive summarization (Cohan et al., 2018; Xu et al., 2020). Coherence relation-aware models have also been shown to generate more coherent texts (Bosselut et al., 2018) within a reinforcement learning setting. Alikhani et al. (2020) extend standard text-based discourse relations to cross-modal scenarios. Our work is the first to leverage coherence relations witin a multimodal dialogue system for safety.
@@
Active learning is a prominent area in machine learning (Settles, 2009), and has gained recent attention for tasks involving natural language (Zhang et al., 2022) such as intent classification, sentence matching, and named entity recognition (Zhang and Zhang, 2019; Bai et al., 2020; Liu et al., 2022). However, active learning has predominantly been focused on classification tasks. Recent works targeting LLMs (Diao et al., 2023; Margatina et al., 2023; Hassan et al., 2024a) also focus on tasks with fixed sets of outputs, leaving active learning for generative tasks largely unexplored (Perlitz et al., 2023; Hassan et al., 2024b). Our work is among the first to apply active learning to generative tasks and multimodal scenarios. Further, we deploy and evaluate active learning in practice, whereas prior work have primarily relied on simulations (Zhang et al., 2022).
@@
With advancement in language models and multimodal models, there has been a surge in research integrating these models with embodied agents (Ahn et al., 2022; Driess et al., 2023; Huang et al., 2022b, a). Most of these recent works focus on extracting plans or executable actions by the robot from pre-trained language models. Different from the aforementioned works, ours is the first to develop a multimodal dialogue system for safety in embodied agents and conduct a real-world user study by deploying it with a robot.
@@ 3.1. Dialogue System
The first input to our dialogue system is an image that contains a potential safety violation. These safety violations may occur during household tasks or within a living environment. For the training phase, images from social media are used, and the learner LLM is fine-tuned with four turns of simulated conversation:
@@
During the user-study phase, the images are captured by an actual robot in household environment and the conversations are not restricted to a specific number of turns, continuing indefinitely.
@@ 3.2. Coherence Relations
Figure 2. Our proposed active learning loop. Informative instances are identified by an external LLM using a composite score. These instances are then distilled using another LLM for discourse relations and responses. These are then used to retrain the learner LLM.
@@ 3.3. Active Learning
We assume there is a pool of unlabeled dataset U but only a subset of labeled data L can be used for training. L is iteratively constructed by querying target output for the most-informative instance. While other active learning scenarios exist (Settles, 2009), we follow the setting of pool-based active learning because of its relevance to our setting, where we can obtain a large number of safety images but obtaining coherence relations and appropriate responses can be challenging.
Active learning for classification tasks typically identifies highly informative instances with measures of uncertainty, such as entropy (Settles, 2009):
@@
In Eq. 1, i ranges over all possible labels.
@@
Standard Active Learning offers label efficiency over random sampling. It can however, induce bias if the model misjudges its confidence (Hassan et al., 2018). Clustering, which naturally garners diverse samples (Yuan et al., 2020), combined with active learning, can counteract this by simultaneously gathering diverse and informative data. We hypothesize that using an LLM on these diverse and informative data would lead to a more representative set of generations. In our clustering-based setting, the unlabeled data is first vectorized and then the vector space is split into $m$ clusters $\{C_{1},C_{2},...C_{m}\}$, where $m$ is a predefined number. Informativeness measure according to Eq. 2 is calculated for each instance within a cluster and most informative samples are chosen from each cluster. These samples are then passed to a distillation LLM for obtaining coherence relation and appropriate response. The combined approach of integrating discourse relation and active learning is summarized in Algorithm 1.
@@ 4.1. Dataset Construction
As the LMM may incur errors in the dataset, the output of the LMM is then annotated by the same graduate annotators. The annotators were asked to take one of three possible actions: i) mark as correct if the LMM output corresponds to a safety violation in the image, ii) if small edits could fix the LMM output, then edit, and iii) discard the safety violation if it does not correspond to the input image. Following this stage of annotation, 107 were discarded, 825 safety violations were retained as is, and 83 were retained after editing, for a total of 908 safety violations retained. Figure 4 shows examples from the dataset. This dataset is the seed dataset $U$ in Algorithm 1.
Figure 3. Distribution of safety-related keywords in our dataset. Safety violations related to mold and stove are more prevalent.
(a) The image shows a knife on the edge of kitchen counter. This can be dangerous as …
@@
Figure 4. Examples of images retrieved from Reddit and their corresponding safety violations, obtained through LMM and manual correction.
@@ 4.2. Dataset Analysis
Table 2. Automated evaluation of dialogue systems. Integrating coherence relations yield higher safety scores, further improved by integration of clustering-based active learning (M-CoDAL). Mistral models are learner models in the active learning paradigm while LLama and Qwen are transfer models fine-tuned with data acquired by active learning.
@@ 5. Experiments
In our experiments, we start with the dataset prepared in Section 4 and proceed with the framework presented in Section 3.
@@ 5.1. Experiment Setup
We use a Mistral 7B (Jiang et al., 2023), a recent and capable Large Language Model to engage in dialogue once a safety violation detected. This is the learner model that is fine-tuned in every iteration of active learning. All fine-tuning is done for 5 epochs with a batch size of 4. We use separate Mistral 7B models to compute composite score for the learner LLM (Eq. 2) and to simulate Turn #2 and Turn #4. The fine-tuned model continues conversation past Turn #4 during the user study.
@@
We use GPT-4o (202, 2023) to distill coherence relations and appropriate responses. GPT-4o is chosen as one of the most capable LLMs and acts as teacher to the smaller open-source learner LLMs. The GPT-4o responses, conditioned by coherence relations, are passed to the learner LLM for fine-tuning.
@@ 5.2. Results
Automated evaluation for our setting is challenging as we observed that standard classifiers such as BERT (Devlin et al., 2019) models trained on datasets such as DiaSafety (Sun et al., 2022b), fail to recognize the nuances necessary for evaluating safety of the dialogue. Thus, for automated evaluation, we deploy a Mistral 7B to determine quality of the generated responses along three dimensions: i) user sentiment score, ii) resolution score, and iii) safety score. The Mistral 7B is prompted to provide a value between 0 and 1.0. In addition, we also calculate average length of the bot response, the user response, as well as number of unique tokens.
@@ 6. User Study
To demonstrate the effectiveness of our proposed system, we conducted a user study that investigated how persuasive and competent a robot powered by M-CoDAL is perceived by users in different safety scenarios. Since people may be more receptive when the safety violation is more severe, such as when there are sharp objects on the ground, we also varied the severity in our study.
@@ 6.1. Study Design
Participants role played two scenarios with different level of severity. In the low-severity scenario, Participants role-played a scenario where they twisted wires and left them on the table. In the high-severity scenario, Participants role-played a scenario where they pretended to cut fruits and vegetables using a knife, then placed the knife at the edge of the table, creating a higher risk of the knife falling off. Figure 5 shows images of these scenarios.
@@
We expect our proposed system M-CoDAL to be more persuasive and competent due to integration of coherence relation and fine-tuning with clustering-based active learning regardless of severity scenario.
@@
We measured the persuasiveness of each robot by asking participants to rate how convincing they found the robot’s suggestions or warnings on a 5-point Likert scale. We also measured the robot’s perceived competence (6 items) and discomfort (6 items) using the Robotic Social Attributes Scale (RoSAS) (Carpinella et al., 2017). At the end of each scenario, we also asked participants which robot they preferred to work with.
@@
Table 3. Findings of the user study in low and high severity scenarios. The robot powered by our proposed system, M-CoDAL, is perceived to be more persuasive, as well as competent compared to the robot powered by GPT-4o.

### v2

@@ Abstract.
When assisting people in daily tasks, robots need to accurately interpret visual cues and respond effectively in diverse safety-critical situations, such as sharp objects on the floor. In this context, we present M-CoDAL, a multimodal-dialogue system specifically designed for embodied agents to better understand and communicate in safety-critical situations. The system leverages discourse coherence relations to enhance its contextual understanding and communication abilities. To train this system, we introduce a novel clustering-based active learning mechanism that utilizes an external Large Language Model (LLM) to identify informative instances. Our approach is evaluated using a newly created multimodal dataset comprising 1K safety violations extracted from 2K Reddit images. These violations are annotated using a Large Multimodal Model (LMM) and verified by human annotators. Results with this dataset demonstrate that our approach improves resolution of safety situations, user sentiment, as well as safety of the conversation. Next, we deploy our dialogue system on a Hello Robot Stretch robot and conduct a within-subject user study with real-world participants. In the study, participants role-play two safety scenarios with different levels of severity with the robot and receive interventions from our model and a baseline system powered by OpenAI’s ChatGPT. The study results corroborate and extend the findings from the automated evaluation, showing that our proposed system is more persuasive in a real-world embodied agent setting.
@@
Embodied Agents; Multimodal Safety; Active Learning; Coherence Theory
@@ 1. Introduction
Embodied agents, such as assistive robots with multimodal capabilities, are poised to become an integral part of our daily lives by helping with household tasks and providing assistance upon request. However, the agent should also play a proactive role and ensure both its and the user’s actions are safe. It needs to not only detect unsafe scenarios but also communicate and persuade the user of the danger. To achieve this, the system needs to understand and adapt to conversational contexts while remaining robust across diverse safety-related scenarios.
To meet this challenge, we integrate theories of coherence relations to enhance contextual understanding and use clustering-based active learning to train a multimodal dialogue system, M-CoDAL, specifically designed for deploying embodied agents for real-world users to provide effective safety advice.
@@
M-CoDAL uses discourse coherence relations to enrich its contextual understanding. Theories of coherence relations originate in understanding and analyzing inferential links to support text interpretation and has subsequently been extended to cross-modal settings (Alikhani et al., 2023). Coherence relations can situate an ongoing scene in the arc of a narrative (Cohn, 2013) or enrich interpretation of communicative actions across modalities (Lascarides and Stone, 2009). For instance, the coherence relation Cause can help interpret why paying attention to a safety violation is important in an image captured by a robot given a certain Condition relation. These relations facilitate dialogue between the human user and the robot. By default, a robot equipped with pretrained models would not explicitly model such inferential links (Alikhani et al., 2023). We hypothesize that embodied agents that parse coherence relations of safety violation in captured images and use them to guide their responses will better understand contexts of safety and engage in safer conversations with humans.
To train M-CoDAL for embodied agents with coherence relation integration, we adopt clustering-based active learning (Settles, 2009). Active learning offers advantage over standard fine-tuning methods by focusing resources toward informative instances. When combined with clustering, active learning has been shown to yield more representative models (Hassan and Alikhani, 2023a). Thus, we hypothesize that applying clustering-based active learning would lead to better coverage of safety scenarios and a safer multimodal dialogue system. Active learning, however, has mostly been addressed in the context of classification tasks and has remained challenging for generative tasks (Perlitz et al., 2023) such as conversational response generation due to difficulty in estimating model uncertainty in a large output space. We make the first inroads in active learning for autonomous agents with conversational capacity by integrating an external LLM that quantifies informativeness of an instance based on a composite score of safety of conversation, resolution, and user sentiment. Further, we distill knowledge from another external LLM to reduce necessary human efforts in active learning. Our work is also one of the first to deploy and evaluate active learning in practice for human-robot interactions.
@@
Finally, we deploy our multimodal dialogue system on a Hello Robot Stretch robot and conduct a within-subject in-person user study where participants interact with our robot in staged unsafety scenarios. 8 participants interacted with the robot operating with GPT-4o or our proposed system, M-CoDAL in low and high severity fake unsafe scenarios. The study results corroborate findings of automated evaluation by showing that participants found our proposed system M-CoDAL to be more persuasive in both conditions. Our qualitative analysis of interviews with the participants also reveals that the participants find M-CoDAL to be more attentive to safety situation and can help the user to be more aware of safety situations. Thus, the key contributions of this paper are:
@@ 2. Related Work
There has been a growing interest in development of multimodal dialogue systems in recent years (Sun et al., 2022a; Lee et al., 2023; Sicilia et al., 2023), especially with the advent of large multimodal models such as GPT-4V (202, 2023) and LLaVA (Liu et al., 2023). Discussions of safety however, has primarily focused on textual domain (Sun et al., 2022b; Weidinger et al., 2022; Dinan et al., 2022; Atwell et al., 2022) through means of text-classification or integrating guardrails within LLMs Dubey et al. (2024). These approaches may not be sufficient to tackle safety situations that may arise in multimodal household scenarios, particularly when the context demands further probing. To our knowledge, our work is the first public work to process visual cues of safety in multimodal dialogues with an embodied agent.
@@
Coherence relations have been proposed as a possible method for controlling generative models and have been shown to aid tasks such as extractive and abstractive summarization (Cohan et al., 2018; Xu et al., 2020). Coherence relation-aware models have also been shown to generate more coherent texts (Bosselut et al., 2018) within a reinforcement learning setting. Alikhani et al. (2020) extend standard text-based discourse relations to cross-modal scenarios. We build on theories of coherence relations to capture the context of safety scenarios and control responses accordingly within a multimodal dialogue system.
@@
Active learning is a prominent area in machine learning (Settles, 2009), and has gained recent attention for tasks involving natural language (Zhang et al., 2022) such as intent classification, sentence matching, and named entity recognition (Zhang and Zhang, 2019; Bai et al., 2020; Liu et al., 2022). However, active learning has predominantly been focused on classification tasks. Recent works targeting LLMs (Diao et al., 2023; Margatina et al., 2023; Hassan et al., 2024a) also focus on tasks with fixed sets of outputs, leaving active learning for generative tasks largely unexplored (Perlitz et al., 2023; Hassan et al., 2024b). Our work is among the first to pioneer active learning for conversational generative tasks and also extend to multimodal scenarios. Further, we deploy and evaluate active learning in practice, whereas prior work have primarily relied on simulations (Zhang et al., 2022).
@@
With advancement in language models and multimodal models, there has been a surge in research to integrate these models with embodied agents (Ahn et al., 2022; Driess et al., 2023; Huang et al., 2022b, a). Most of these recent works focus on extracting plans or executable actions by the robot from pre-trained language models. In contrast, our work enables the robot to observe safety scenarios in its surroundings and communicate with the user. Our real-world user study, conducted by enacting scenarios letting the users interact with the robot, also reveals insights on how human users may perceive such an agent.
@@ 3.1. Dialogue System
The first input to our dialogue system is an image that contains a potential safety violation. These safety violations may occur during household tasks or within a living environment. For training, the learner LLM is fine-tuned with four turns of simulated conversation:
@@
Turn #1 is obtained by processing image in our dataset and the subsequent turns are simulated by LLMs for training. During the user-study phase, the images are captured by an actual robot in household environment and the conversations are not restricted to a specific number of turns, continuing indefinitely.
@@ 3.2. Coherence Relations
Figure 2. In our active learning loop, informative instances are identified by an external LLM using a composite score. These are distilled using another LLM for coherence relations and responses and are used to retrain the learner LLM.
@@ 3.3. Active Learning
We assume there is a pool of unlabeled dataset U but only a subset of labeled data L can be used for training. L is iteratively constructed by querying target output for the most-informative instance. While other active learning scenarios exist (Settles, 2009), we follow the setting of pool-based active learning because of its relevance to our setting, where we can obtain a large number of safety images but obtaining coherence relations and appropriate responses can be challenging. In a standard classification setting, active learning would typically identify informative instances from this pool with measures of uncertainty, such as entropy (Settles, 2009):
@@
In Eq. 1, $y_{i}$ is the $i^{th}$ possible output for input $x$.
@@
Standard Active Learning offers label efficiency over random sampling. It can however, induce bias if the model misjudges its confidence (Hassan et al., 2018). Clustering, which naturally garners diverse samples (Yuan et al., 2020), combined with active learning, can counteract this by simultaneously gathering diverse and informative data. We hypothesize that using an LLM on these diverse and informative data would lead to a more representative set of generations. In our clustering-based setting, the unlabeled data is first vectorized and then the vector space is split into $m$ clusters $\{C_{1},C_{2},...C_{m}\}$, where $m$ is a predefined number. Informativeness measure according to Eq. 2 is calculated for each instance within a cluster and most informative samples are chosen from each cluster. These samples are then passed to a distillation LLM for obtaining coherence relation and appropriate response. The combined approach of integrating discourse relation and active learning is illustrated in Figure 2 and summarized in Algorithm 1.
@@ 4.1. Dataset Construction
As the LMM may incur errors in the dataset, the output of the LMM is then annotated by the same graduate annotators. The annotators were asked to take one of three possible actions: i) mark as correct if the LMM output corresponds to a safety violation in the image, ii) if small edits could fix the LMM output, then edit, and iii) discard the safety violation if it does not correspond to the input image. Following this stage of annotation, 107 were discarded, 825 safety violations were retained as is, and 83 were retained after editing, for a total of 908 safety violations. Figure 4 shows examples from the dataset. This dataset is the seed dataset $U$ in Algorithm 1.
Figure 3. Distribution of safety keywords in our dataset. Safety violations with mold and stove are more prevalent.
(a) The image shows a knife on the edge of kitchen counter. This can be dangerous …
@@
Figure 4. Examples of images retrieved from Reddit and their safety violations, obtained through LMM and manual correction.
@@ 4.2. Dataset Analysis
Table 2. Automated evaluation of dialogue systems. Integrating coherence relations yield higher safety scores, further improved by clustering-based active learning (M-CoDAL). Mistral models are learner models while LLama and Qwen are transfer models.
@@ 5.1. Experiment Setup
We use a Mistral 7B (Jiang et al., 2023), a recent and capable Large Language Model to engage in dialogue once a safety violation is detected. This is the learner model that is fine-tuned in every iteration of active learning. All fine-tuning is done for 5 epochs with a batch size of 4. We use separate Mistral 7B models to compute composite scores for the learner LLM (Eq. 2) and to simulate Turn #2 and Turn #4. The fine-tuned model continues the conversation past Turn #4 during the user study.
@@
We use GPT-4o (202, 2023) to distill coherence relations and appropriate responses. GPT-4o is chosen as one of the most capable LLMs and acts as a teacher to the smaller open-source learner LLMs. The GPT-4o responses, conditioned by coherence relations, are passed to the learner LLM for fine-tuning.
@@ 5.2. Results
Automated evaluation for our setting is challenging as we observed that standard classifiers such as BERT (Devlin et al., 2019) models trained on datasets such as DiaSafety (Sun et al., 2022b), fail to recognize the nuances necessary for evaluating safety from the dialogue. Thus, for automated evaluation, we deploy an external Mistral 7B to determine the quality of the generated responses along three dimensions: i) user sentiment score, ii) resolution score, and iii) safety score. The Mistral 7B is prompted to provide a value between 0 and 1.0. In addition, we also calculate the average length of the bot response, the user response, as well as the number of unique tokens.
@@ 6. User Study
To demonstrate the effectiveness of our proposed system, we conducted a user study that investigated how persuasive and competent a robot powered by M-CoDAL (Mistral variation) is perceived by users in different safety scenarios. Since users may be more receptive when the safety violation is more severe, such as when there are sharp objects on the ground, we varied the severity in our study.
@@ 6.1. Study Design
Participants role played two scenarios with different levels of severity. In the low-severity scenario, Participants role-played a scenario where they twisted wires and left them on the table. In the high-severity scenario, Participants role-played a scenario where they pretended to cut fruits and vegetables using a knife, then placed the knife at the edge of the table, creating a higher risk of the knife falling off. Fake tools and appliances were used so that no real unsafe scenario would occur for the participants. Figure 5 shows images of these scenarios.
@@
We expect our proposed system M-CoDAL to be more persuasive and competent due to the integration of coherence relation and fine-tuning with clustering-based active learning.
@@
We measured the persuasiveness of each robot by asking participants to rate how convincing they found the robot’s suggestions or warnings on a 5-point Likert scale. We also measured the robot’s perceived competence (6 items) and discomfort (6 items) using the Robotic Social Attributes Scale (RoSAS) (Carpinella et al., 2017). We also asked the participants which robot they preferred to work with.
Table 3. Findings of the user study in low and high severity scenarios. The robot powered by our proposed system, M-CoDAL, is perceived to be more persuasive compared to the robot powered by GPT-4o.